OpenStudy (anonymous):

What is the slope of line LM given L(9, -2) and M(3, -5)?

OpenStudy (anonymous):

Okay so the slope is the gradient. Gradient is given by (y1 - y2)/(x1-x2) where y1 stands for the first group of coordinates which can be either L(9, -2) or M(3, -5), it doesn't matter which you decide is the first group as the answer will be the same. So then gradient equals (-2 - -5)/(9 - 3) = 3/6 = 1/2 Just make sure to use take both x1 and y1 from the SAME coordinates: L or M (it doesn't matter which but if y1 is the y-value from M then x1 has to be the x-value from M as well).